the graph of the equation y=x cubed -x is symmetric with respect to which of the following?
A)the x-axis
B)the y-axis
C)the origin
D)none of these

Answer by stanbon(75887)

the graph of the equation y=x cubed -x is symmetric with respect to which of the following?
A)the x-axis
B)the y-axis
C)the origin
D)none of these
-----------------------------
Look at f(x), f(-x), and -f(-x)
f(x) = x^3-x
f(-x) = -(x^3)-x
-f(-x) = x^3 + x
--------------------------
Since f(x) = -f(-x) the function is symmetric to the origin.
====================================================================
Cheers,
Stan H.